three consecutive integers where the square of the smaller number is 4 more than 8 time the largest what are the three intergers?
Answer by JulietG(1812)

Let A be the largest number and C the smallest.
A = B+1, or C+2 (because they are consecutive)
B = C+1, or A-1
C = A-2, or B-1
c^2 - 4 = 8A [the square of the smaller number is 4 more than 8 time the largest]
Substitute the known value of A into bottom equation.
c^2 - 4 = 8(C+2)
Distribute
c^2 - 4 = 8C + 16
Add 4 to each side
c^2 = 8C + 20
c = 10
.
If C is 10, then B is 11, and A is 12.
The square of C is 100, which is 4 more than 8*12